OpenStudy (heart_offire):

A line has a slope of 1/2 and passes through the point (0,5). What is the equation of the line? A) y = 1/2x + 5 B) y = 1/2x - 5 C) y = 5x + 1/2 D) y = -5x + 1/2 i choose A @pooja195

pooja195 (pooja195):

That's correct well done! :)

pooja195 (pooja195):

y=mx+b m=slope b=y intercept it is defined: y=(0,5) <---5 m=1/2 y=1/2x+5 Just to explain ;)